OpenStudy (anonymous):

The table below shows the number of marbles of different colors in a bag: Ursula draws a marble from the bag randomly without looking. She then draws another marble from the bag without replacing the first one. Which expression shows the probability of drawing black marbles in both the trials?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

B AND C

OpenStudy (anonymous):

@welshfella @Vocaloid @vera_ewing @omgitsjc @LegendarySadist

OpenStudy (anonymous):

@Kash_TheSmartGuy

OpenStudy (anonymous):

Where is the table?

OpenStudy (anonymous):

here

OpenStudy (anonymous):

i was thinking the answer would be a but idk

OpenStudy (anonymous):

Are you sure that's the table? The question asks about black marbles, but the table you provided only talks about Red, blue, green, and purple.

OpenStudy (anonymous):

crap wrong table

OpenStudy (anonymous):

there but i was thinking (a) but im not sure

OpenStudy (anonymous):

When you look for a probability of one thing it's simply \[\large \sf \frac{Desired~object}{Total~objcets}\] Which in this case is \[\large \sf \frac{black~marbles}{Total~marbles}\] When you look for it to happen twice, you multiply the probabilities. Since we're assuming the first will end up with a loss of one black ( and one total ) the second probability would look like this \[\large \sf \frac{black~marbles-1}{Total~marbles-1}\] Now you would just multiply the probabilities together so \[\large \sf \frac{black~marbles}{Total~marbles}~\times ~ \large \sf \frac{black~marbles-1}{Total~marbles-1}\]

OpenStudy (anonymous):

would it be a thats what i got

OpenStudy (anonymous):

No, it's not a

OpenStudy (anonymous):

c

OpenStudy (anonymous):

or b idk

OpenStudy (anonymous):

Why not just plug the numbers into what I gave you?

OpenStudy (anonymous):

\[\large \sf \frac{black~marbles}{Total~marbles}~\times ~ \large \sf \frac{black~marbles-1}{Total~marbles-1}\]

OpenStudy (anonymous):

idkhowwwwwwwwwwwwwwwwwwwwwwwwwwwwww

OpenStudy (anonymous):

How many black marbles are there?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

How many total marbles are there?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

Now plug 10 in for black marbles and 16 for total marbles \[\large \sf \frac{black~marbles}{Total~marbles}~\times ~ \large \sf \frac{black~marbles-1}{Total~marbles-1}\]

OpenStudy (anonymous):

d

OpenStudy (anonymous):

You only showed A B and C as your options.

OpenStudy (anonymous):

soooooooo what is it @LegendarySadist